# Calibrating Bayesian Learning via Regularization, Confidence Minimization, and Selective Inference

### v2

@@ Abstract
The application of artificial intelligence (AI) models in fields such as engineering is limited by the known difficulty of quantifying the reliability of an AI’s decision. A well-calibrated AI model must correctly report its accuracy on in-distribution (ID) inputs, while also enabling the detection of out-of-distribution (OOD) inputs. A conventional approach to improve calibration is the application of Bayesian ensembling. However, owing to computational limitations and model misspecification, practical ensembling strategies do not necessarily enhance calibration. This paper proposes an extension of variational inference (VI)-based Bayesian learning that integrates calibration regularization for improved ID performance, confidence minimization for OOD detection, and selective calibration to ensure a synergistic use of calibration regularization and confidence minimization. The scheme is constructed successively by first introducing calibration-regularized Bayesian learning (CBNN), then incorporating out-of-distribution confidence minimization (OCM) to yield CBNN-OCM, and finally integrating also selective calibration to produce selective CBNN-OCM (SCBNN-OCM). Selective calibration rejects inputs for which the calibration performance is expected to be insufficient. Numerical results illustrate the trade-offs between ID accuracy, ID calibration, and OOD calibration attained by both frequentist and Bayesian learning methods. Among the main conclusions, SCBNN-OCM is seen to achieve best ID and OOD performance as compared to existing state-of-the-art approaches at the cost of rejecting a sufficiently large number of inputs.
Index Terms:
Bayesian learning, calibration, OOD detection, selective calibration
@@ I-A Context and Motivation
TABLE I: Comparison of most related works
reference $\text{learning}^{*}$ calibration regularization OOD confidence minimization selective inference
[4], [5, 6, 7, 8] F
@@
phenomenon of the “hallucinations” of large language models [13, 14, 15, 16]. Indeed, in safety-critical applications, calibration – i.e., the property of a model to know when it does not know – is arguably just as important as accuracy [17, 18, 19, 20].
@@
A conventional approach to improve ID calibration is the application of Bayesian ensembling. A Bayesian neural network (BNN) captures epistemic uncertainty via a distribution over the weights of the model, making it possible to evaluate reliability via the level of consensus between models drawn from the model distribution [2, 21, 9, 22]. However, owing to computational limitations and model misspecification, practical ensembling strategies do not necessarily enhance ID calibration [23, 24, 25, 26, 27]. Furthermore, improvements in ID calibration may not necessarily improve OOD detection [2, 28, 29].
This paper proposes an extension of variational inference (VI)-based Bayesian learning that integrates calibration regularization for improved ID performance [4], confidence minimization for enhanced OOD detection [11], and selective calibration to ensure a synergistic use of calibration regularization and confidence minimization [12]. Prior art only considered the separate applications of these techniques to conventional frequentist neural networks (FNNs), and, to the best of our knowledge, no study has been done on integrating the three approaches and on adapting them for use in the training of BNNs.
Figure 1: (a) Standard frequentist neural networks (FNNs) generally fail to provide well-calibrated decisions, and improved in-distribution (ID) calibration can be achieved via Bayesian neural networks (BNNs) [10]. (b) Calibration regularization improves ID calibration via a regularizer that penalizes calibration errors [4]. (c) Out-of-distribution confidence minimization (OCM) injects OOD examples during training to improve OOD detection [11]. (d) Selective calibration further improves both ID and OOD calibration by only producing decisions for inputs at which uncertainty quantification is deemed to be sufficiently reliable. Prior works [4, 11, 12] introduced calibration regularization, OCM, and selective calibration as separate methods for FNNs, while this work presents an integrated approach encompassing all three solutions for BNNs.
@@ I-B State of the Art
As mentioned, improving ID calibration does not necessarily enhance OOD detection, and tailored methods are required for this purpose [11, 33, 2, 21, 9, 32]. As illustrated in Fig. 1(c), out-of-distribution confidence minimization (OCM) augments the data set with OOD inputs, which are used to specify a regularizer that favors high uncertainty outside the training set [11]. However, in so doing, OCM can hurt ID calibration, making the model excessively conservative.
Overall, there is generally a trade-off between OOD and ID calibration, and schemes designed to optimize either criterion may hurt the other [11]. A possible solution to this problem is to properly select inputs on which to make a decision. Selective inference techniques typically focus on enhancing the accuracy of the model by selecting inputs with the largest model confidence among the selected inputs [34, 35, 34, 36, 37, 12]. As a result, these techniques may end up exacerbating issues with ID calibration by favoring over-confidence. In contrast, as shown in Fig. 1(d), selective calibration learns how to select inputs that are expected to have a low gap between accuracy and confidence, thus boosting ID calibration [12].
All the papers presented so far have focused on improving ID and OOD calibration for FNNs, and are thus limited in the capacity to provide reliable decisions that fully account for epistemic uncertainty.
@@ I-C Main Contributions
This paper proposes an extension of VI-based BNN learning that integrates calibration regularization for improved ID performance, OCM for OOD detection, and selective calibration to ensure a synergistic use of calibration regularization and OCM. As illustrated in Fig. 1, the scheme is constructed successively by first introducing calibration-regularized Bayesian learning (CBNN), then incorporating OCM to yield CBNN-OCM, and finally integrating also selective calibration to produce selective CBNN-OCM (SCBNN-OCM).
@@
  • We propose CBNN, a novel variant of VI-based BNNs that improve the ID calibration performance in the presence of computational complexity constraints and model misspecification by adding a calibration-aware regularizer (see Fig. 1(b)).
  • In order to improve OOD detection, CBNN is extended via OCM, yielding CBNN-OCM, by incorporating an additional regularizer that penalizes confidence on OOD data (see Fig. 1(c)).
@@
  • Extensive experimental results on real-world image classification task, including CIFAR-100 data set [38] and TinyImageNet data set [39], illustrate the trade-offs between ID accuracy, ID calibration, and OOD calibration attained by both FNN and BNN. Among the main conclusions, SCBNN-OCM is seen to achieve best ID and OOD performance as compared to existing state-of-the-art approaches at the cost of rejecting a sufficiently large number of inputs.
A table summarizing the comparison of our paper to existing studies can be found in Table LABEL:tab:my_label. In this regard, it is noted that versions of CBNN have been presented in [10, 9]. The authors in [9] proposed a new calibration-aware regularization term for Bayesian learning, AvUC, whose limitations as a calibration-aware regularizer were demonstrated in [8]. Moreover, reference [10] considered a more general CBNN framework, while focusing solely on ID calibration performance.
@@ II-A Problem Definition: In-Distribution Calibration
The extent to which condition (3) is satisfied is typically evaluated via the reliability diagram [41] and the expected calibration error (ECE) [42, 30]. Both approaches estimate the probability in (3) using a test data set $\mathcal{D}^{\text{te}}=\{(x_{i}^{\text{te}},y_{i}^{\text{te}})\}_{i=1}^{|\mathcal{D}^{\text{te}}|}$, with each pair $(x_{i}^{\text{te}},y_{i}^{\text{te}})\sim p(x,y)$ being drawn independently of the training data set $\mathcal{D}^{\text{tr}}$.
@@
[…]
@@
Each $m$-th bin is associated with a per-bin confidence $\mathrm{conf}(\mathcal{B}_{m})={1}/{|\mathcal{B}_{m}|}\sum_{i\in\mathcal{B}_{m}}r_{i}^{\text{te}}$, and with a per-bin accuracy $\mathrm{acc}(\mathcal{B}_{m})={1}/{|\mathcal{B}_{m}|}\sum_{i\in\mathcal{B}_{m}}c_{i}^{\text{te}}$. An estimate of the reliability diagram is now obtained by plotting $\mathrm{acc}(\mathcal{B}_{m})$ as a function of $\mathrm{conf}(\mathcal{B}_{m})$ across all bins $m=1,...,M$.
@@ II-C Background: Bayesian Learning
In practice, the optimization (13) over distribution $q(\theta)$ is often carried out over a parameterized class of distribution via variational inference (VI) [43, 44] (see [22, Ch. 12] for an overview). Specifically, denoting as $q(\theta|\varphi)$ a distribution over parameter vector $\theta$ parameterized by a vector $\varphi$, the problem (13) is simplified as
@@ IV-A Problem Definition: Selective Classification
By adopting the design criterion (27), selective classification aims at choosing only the inputs $x$ for pairs $(x,y)$ with large confidence levels $p(y|x,\mathcal{D}^{\text{tr}})$. As anticipated, this objective is sensible if the underlying model $p(y|x,\mathcal{D}^{\text{tr}})$ is sufficiently well calibrated. Reference [34] addressed the constrained optimization in (28) by converting it into an unconstrained problem via a quadratic penalty function (see [34, Eq. (2)]).
@@ IV-B1 Training the Selector
where ~$\tilde{\mathcal{E}}^{\text{S-Cal}}(\phi|\mathcal{D}^{\text{val}})$, defined in (34),
~$\tilde{\mathcal{E}}^{\text{S-Cal}}(\phi|\mathcal{D}^{\text{val}})=\left(\sum_{i=1}^{|\mathcal{D}^{\text{val}}|}\sum_{j=1}^{|\mathcal{D}^{\text{val}}|}(c_{i}-{r}_{i})(c_{j}-{r}_{j})\tilde{g}(r_{i},s_{i}|\phi)\tilde{g}(r_{j},s_{j}|\phi)\kappa({r}_{i},{r}_{j})\right)^{\frac{1}{2}},$ (34)
uses confidence score $r_{i}=r(x_{i}^{\text{val}}|\theta^{\text{FNN}})$ and non-parametric outlier score vector $s_{i}=s(x_{i}^{\text{val}}|\theta^{\text{FNN}})$, with a coefficient $\eta\geq 0$ that balances selective calibration performance and coverage. Note that function (34) is the numerator of (30) with the relaxed selector $\tilde{g}(r(x|\theta^{\text{FNN}}),s(x|\theta^{\text{FNN}})|\phi)$ in lieu of $g(x|\phi)$, and that $-\log(\cdot)$ plays the role of a barrier function to enforce the coverage rate constraint [52].
@@ IV-B2 Non-Parametric Outlier Score Vector
$s_{1}(x|\theta^{\text{FNN}})=\frac{1}{|\mathcal{X}^{\text{tr}}|}\sum_{i=1}^{|\mathcal{X}^{\text{tr}}|}\kappa\left(||z-z^{\text{tr}}_{i}||\right),$ (35)
where $\kappa(\cdot)$ is a kernel function; $z$ is the output of the last hidden layer of the parametrized network $p(y|x,\theta^{\text{FNN}})$; and, similarly, $z_{i}^{\text{tr}}$ is the corresponding feature vector for $p(y|x_{i},\theta^{\text{FNN}})$. Details regarding the other elements of the score vector $s(x|\theta^{\text{FNN}})$ can be found in Appendix -A.
@@ IV-C Proposed Method: Bayesian Learning with Selective Calibration
$\phi^{\text{SCBNN-OCM}}=\arg\min_{\phi}\Bigg{\{}\tilde{\mathcal{E}}^{\text{S-Cal}}(\phi|\mathcal{D}^{\text{val}})$
$-\eta\cdot\sum_{x\in\mathcal{D}^{\text{val}}}\log\tilde{g}(\bar{r}(x|\varphi^{\text{CBNN-OCM}}),\bar{s}(x|\varphi^{\text{CBNN-OCM}})|\phi)\Bigg{\}},$ (39)
where ~$\tilde{\mathcal{E}}^{\text{S-Cal}}(\phi|\mathcal{D}^{\text{val}})$, defined in (LABEL:eq:sel-cal-loss-bayes),
~$\tilde{\mathcal{E}}^{\text{S-Cal}}(\phi|\mathcal{D}^{\text{val}})=\left(\sum_{i=1}^{|\mathcal{D}^{\text{val}}|}\sum_{j=1}^{|\mathcal{D}^{\text{val}}|}(c_{i}-{\bar{r}}_{i})(c_{j}-{\bar{r}}_{j})\tilde{g}(\bar{r}_{i},\bar{s}_{i}|\phi)\tilde{g}(\bar{r}_{j},\bar{s}_{j}|\phi)\kappa({\bar{r}}_{i},{\bar{r}}_{j})\right)^{\frac{1}{2}},$ (40)
@@
$g(x|\phi^{\text{SCBNN-OCM}})$
$=\mathbbm{1}\left(\tilde{g}(\bar{r}(x|\varphi^{\text{CBNN-OCM}}),\bar{s}(x|\varphi^{\text{CBNN-OCM}})|\phi^{\text{SCBNN-OCM}})\geq\tau\right),$ (41)

### v3

@@ Abstract
The application of artificial intelligence (AI) models in fields such as engineering is limited by the known difficulty of quantifying the reliability of an AI’s decision. A well-calibrated AI model must correctly report its accuracy on in-distribution (ID) inputs, while also enabling the detection of out-of-distribution (OOD) inputs. Conventional solutions are effective at addressing only one of these two conflicting requirements. This paper proposes a novel solution that endows Bayesian learning with enhanced ID calibration and OOD detection capabilities by integrating calibration regularization for improved ID performance, confidence minimization for OOD detection, and selective calibration to ensure a synergistic use of calibration regularization and confidence minimization. Selective calibration rejects inputs for which the calibration performance is expected to be insufficient, supporting effective OOD detection, while also ensuring ID calibration. Prior art had only considered these ideas in isolation and for frequentist learning. Numerical results illustrate the trade-offs between ID accuracy, ID calibration, and OOD calibration, showing that the proposed novel Bayesian approach achieves the best ID and OOD performance compared to existing state-of-the-art approaches, at the cost of rejecting a fraction of the inputs.
Abstract
This supplementary document consists of four sections. Ablation studies on ID calibration, OOD detection, and selective calibration are conducted to illustrate the impact of hyperparameters, i.e., weight of regularize, in Sec. VII, Sec. VIII, and Sec. IX, respectively. In Sec. X, to showcase the general performance gains of the proposed framework, we present the additional evaluations on a different backbone. We refer to appendix for all the experimental details.
@@ I-A Context and Motivation
TABLE I: Comparison with state of the art
reference $\text{learning}^{*}$
calibration
regularization
OOD
confidence
minimization
selective
inference
[4],
[5, 6, 7, 8]
F
@@
* “F” == frequentist learning; “B” == Bayesian learning.
phenomenon of the “hallucinations” of large language models [13, 14, 15]. Indeed, in safety-critical applications, calibration – i.e., the property of a model to know when it does not know – is arguably just as important as accuracy [16, 17, 18].
@@
A conventional approach to improve ID calibration is the application of Bayesian ensembling. A Bayesian neural network (BNN) captures epistemic uncertainty via a distribution over the weights of the model, making it possible to evaluate reliability via the level of consensus between models drawn from the model distribution [2, 9, 19]. However, owing to computational limitations and model misspecification, practical ensembling strategies do not necessarily enhance ID calibration [20, 21, 22]. Furthermore, improvements in ID calibration often degrade OOD detection performance [2, 23, 24].
Focusing on conventional frequentist neural networks (FNNs), prior art has introduced a number of notable methods to separately enhance ID calibration and OOD detection. In particular, first, calibration regularization improves ID performance by penalizing excessively confident prediction [4]. Second, confidence minimization enhances OOD detection by exposing the model to OOD samples during training [11]. Finally, selective calibration adaptively decides to reject inputs for which the calibration error is estimated to be excessively large [12]. This work proposes a novel training strategy that endows BNNs with enhanced ID calibration and OOD detection capabilities by integrating calibration regularization for improved ID performance, confidence minimization for OOD detection, and selective calibration to ensure a synergistic use of calibration regularization and confidence minimization. A table summarizing the comparison of our paper to existing studies can be found in Table I.
Figure 1: (a) Standard frequentist neural networks (FNNs) generally fail to provide well-calibrated decisions, and improved in-distribution (ID) calibration can be achieved via Bayesian neural networks (BNNs) [10]. (b) Calibration regularization improves ID calibration via a regularizer that penalizes calibration errors [4]. (c) Out-of-distribution confidence minimization (OCM) injects OOD examples during training to improve OOD detection [11]. (d) Selective calibration further improves both ID and OOD calibration by only producing decisions for inputs at which uncertainty quantification is deemed to be sufficiently reliable. Prior works [4, 11, 12] introduced calibration regularization, OCM, and selective calibration as separate methods for FNNs. In contrast, this work presents an integrated training method for BNNs that integrates calibration regularization for improved ID performance, confidence minimization for OOD detection, and selective calibration to ensure a synergistic use of calibration regularization and confidence minimization.
@@ I-B State of the Art
As mentioned, improving ID calibration often degrades OOD detection performance, and tailored methods are required to enhance OOD performance [11, 28, 2, 29, 9, 27]. As illustrated in Fig. 1(c), out-of-distribution confidence minimization (OCM) augments the data set with OOD inputs, which are used to specify a regularizer that favors high uncertainty outside the training set [11]. However, in so doing, OCM can hurt ID calibration, making the model excessively conservative.
Overall, there is generally a trade-off between OOD and ID calibration, and schemes designed to optimize either criterion may hurt the other [11].
Selective inference techniques typically focus on enhancing the accuracy of the model by selecting inputs with the largest model confidence among the selected inputs [30, 31, 32, 12]. As a result, these techniques may end up exacerbating issues with ID calibration by favoring over-confidence. In contrast, as shown in Fig. 1(d), selective calibration learns how to select inputs that are expected to have a low gap between accuracy and confidence, thus boosting ID calibration [12].
All the papers presented so far have focused on improving either ID or OOD calibration. Furthermore, by focusing on FNNs, these methods are limited in the capacity to provide reliable decisions that fully account for epistemic uncertainty.
@@ I-C Main Contributions
This paper proposes a novel methodology for BNN training that integrates calibration regularization for improved ID performance, OCM for OOD detection, and selective calibration to ensure a synergistic use of calibration regularization and OCM. As illustrated in Fig. 1, this paper constructs this scheme successively by first introducing calibration-regularized Bayesian learning (CBNN); then incorporating OCM to yield CBNN-OCM; and finally integrating also selective calibration to produce the proposed selective CBNN-OCM (SCBNN-OCM).
@@
  • We propose CBNN, a novel BNN training scheme that improves the ID calibration performance in the presence of computational complexity constraints and model misspecification by adding a calibration-aware regularizer (see Fig. 1(b)).
  • In order to improve OOD detection, CBNN is extended by incorporating an additional regularizer that penalizes confidence on OOD data via OCM. The resulting scheme is referred to as CBNN-OCM (see Fig. 1(c)).
@@
  • Extensive experimental results on real-world image classification task, including CIFAR-100 data set [33] and TinyImageNet data set [34], illustrate the trade-offs between ID accuracy, ID calibration, and OOD calibration attained by both FNN and BNN. Among the main conclusions, SCBNN-OCM is seen to achieve best ID and OOD performance as compared to existing state-of-the-art approaches as long as a fraction of inputs is rejected.
Versions of CBNN have appeared in [10, 9], with [10] being an earlier conference version of this work. The authors in [9] proposed a new calibration-aware regularization term for Bayesian learning, AvUC, whose limitations as a calibration-aware regularizer were demonstrated in [8]. Reference [10] studied a more general CBNN framework compared to [9], while focusing solely on ID calibration performance.
@@ II-A Problem Definition: In-Distribution Calibration
The extent to which condition (3) is satisfied is typically evaluated via the reliability diagram [36] and the expected calibration error (ECE) [25]. Both approaches estimate the probability in (3) using a test data set $\mathcal{D}^{\text{te}}=\{(x_{i}^{\text{te}},y_{i}^{\text{te}})\}_{i=1}^{|\mathcal{D}^{\text{te}}|}$, with each pair $(x_{i}^{\text{te}},y_{i}^{\text{te}})\sim p(x,y)$ being drawn independently of the training data set $\mathcal{D}^{\text{tr}}$.
@@
with the indicator function $\mathbbm{1}(\cdot)$ defined as $\mathbbm{1}(\text{true})=1$ and $\mathbbm{1}(\text{false})=0$. To enable an estimate of the reliability diagram, we partition the test data points into bins with approximately the same confidence level (4). Specifically, the $m$-th bin $\mathcal{B}_{m}$ contains indices $i$ of the test inputs that have confidence scores lying in the interval $(\frac{m-1}{M},\frac{m}{M}]$, i.e., $\mathcal{B}_{m}=\{i\in\{1,...,M\}:r_{i}^{\text{te}}\in(\frac{m-1}{M},\frac{m}{M}]\}$.
@@
Each $m$-th bin is associated with a per-bin confidence $\mathrm{conf}(\mathcal{B}_{m})={1}/{|\mathcal{B}_{m}|}\sum_{i\in\mathcal{B}_{m}}r_{i}^{\text{te}}$, and with a per-bin accuracy $\mathrm{acc}(\mathcal{B}_{m})={1}/{|\mathcal{B}_{m}|}\sum_{i\in\mathcal{B}_{m}}c_{i}^{\text{te}}$. An estimate of the reliability diagram is now obtained by plotting $\mathrm{acc}(\mathcal{B}_{m})$ as a function of $\mathrm{conf}(\mathcal{B}_{m})$ across all bins m=1,,Mm=1,...,M.
@@ II-C Background: Bayesian Learning
In practice, the optimization (13) over distribution $q(\theta)$ is often carried out over a parameterized class of distribution via variational inference (VI) [38, 39] (see [19, Ch. 12] for an overview). Specifically, denoting as $q(\theta|\varphi)$ a distribution over parameter vector $\theta$ parameterized by a vector $\varphi$, the problem (13) is simplified as
@@ IV-A Problem Definition: Selective Classification
By adopting the design criterion (27), selective classification aims at choosing only the inputs $x$ for pairs $(x,y)$ with large confidence levels $p(y|x,\mathcal{D}^{\text{tr}})$. As anticipated, this objective is sensible if the underlying model $p(y|x,\mathcal{D}^{\text{tr}})$ is sufficiently well calibrated. Reference [30] addressed the constrained optimization in (28) by converting it into an unconstrained problem via a quadratic penalty function (see [30, Eq. (2)]).
@@ IV-B1 Training the Selector
where $\tilde{\mathcal{E}}^{\text{S-Cal}}(\phi|\mathcal{D}^{\text{val}})$, defined in (34),
$\tilde{\mathcal{E}}^{\text{S-Cal}}(\phi|\mathcal{D}^{\text{val}})=\left(\sum_{i=1}^{|\mathcal{D}^{\text{val}}|}\sum_{j=1}^{|\mathcal{D}^{\text{val}}|}(c_{i}-{r}_{i})(c_{j}-{r}_{j})\tilde{g}(r_{i},s_{i}|\phi)\tilde{g}(r_{j},s_{j}|\phi)\kappa({r}_{i},{r}_{j})\right)^{\frac{1}{2}},$ (34)
uses confidence score $r_{i}=r(x_{i}^{\text{val}}|\theta^{\text{FNN}})$ and non-parametric outlier score vector $s_{i}=s(x_{i}^{\text{val}}|\theta^{\text{FNN}})$, with a coefficient $\eta\geq 0$ that balances selective calibration performance and coverage. Note that function (34) is the numerator of (30) with the relaxed selector $\tilde{g}(r(x|\theta^{\text{FNN}}),s(x|\theta^{\text{FNN}})|\phi)$ in lieu of $g(x|\phi)$, and that $\log(\cdot)$ plays the role of a barrier function to enforce the coverage rate constraint [49].
@@ IV-B2 Non-Parametric Outlier Score Vector
[…]
where $\kappa(\cdot)$ is a kernel function; $z$ is the output of the last hidden layer of the parametrized network $p(y|x,\theta^{\text{FNN}})$; and, similarly, $z_{i}^{\text{tr}}$ is the corresponding feature vector for $p(y|x_{i},\theta^{\text{FNN}})$. Details regarding the other elements of the score vector $s(x|\theta^{\text{FNN}})$ can be found in Appendix in the supplementary material.
@@ IV-C Proposed Method: Bayesian Learning with Selective Calibration
$\phi^{\text{SCBNN-OCM}}=\arg\min_{\phi}\Bigg\{\tilde{\mathcal{E}}^{\text{S-Cal}}(\phi|\mathcal{D}^{\text{val}})$ (39)
$-\eta\cdot\sum_{x\in\mathcal{D}^{\text{val}}}\log\tilde{g}(\bar{r}(x|\varphi^{\text{CBNN-OCM}}),\bar{s}(x|\varphi^{\text{CBNN-OCM}})|\phi)\Bigg\},$
where $\tilde{\mathcal{E}}^{\text{S-Cal}}(\phi|\mathcal{D}^{\text{val}})$, defined in (LABEL:eq:sel-cal-loss-bayes),
$\tilde{\mathcal{E}}^{\text{S-Cal}}(\phi|\mathcal{D}^{\text{val}})=\left(\sum_{i=1}^{|\mathcal{D}^{\text{val}}|}\sum_{j=1}^{|\mathcal{D}^{\text{val}}|}(c_{i}-{\bar{r}}_{i})(c_{j}-{\bar{r}}_{j})\tilde{g}(\bar{r}_{i},\bar{s}_{i}|\phi)\tilde{g}(\bar{r}_{j},\bar{s}_{j}|\phi)\kappa({\bar{r}}_{i},{\bar{r}}_{j})\right)^{\frac{1}{2}},$ (40)
@@
$g(x|\phi^{\text{SCBNN-OCM}})$ (41)
$=\mathbbm{1}\left(\tilde{g}(\bar{r}(x|\varphi^{\text{CBNN-OCM}}),\bar{s}(x|\varphi^{\text{CBNN-OCM}})|\phi^{\text{SCBNN-OCM}})\geq\tau\right),$